OpenStudy (anonymous):

Find the maximum and minimum values of the function on the given interval. y = x^2 – 2x – 7, [-2, 0]

myininaya (myininaya):

Find y' have you?

myininaya (myininaya):

And actually we don't need calculus for this one

myininaya (myininaya):

This is a parabola you can put this in vertex form

myininaya (myininaya):

to determine the vertex which will be a min

myininaya (myininaya):

since the parabola is concave up

myininaya (myininaya):

But we do have endpoints

myininaya (myininaya):

so we need to see if the x-coordinate of the vertex is btw -2 and 0

myininaya (myininaya):

Which way do you prefer? Cal or Algebra?

OpenStudy (anonymous):

actually am doing a calculus based course,so i would like to know it in calculus :)

myininaya (myininaya):

SO have you found y'?

OpenStudy (anonymous):

that would simply be 2x-2

myininaya (myininaya):

\[y'=(x^2)'-(2x)'-(7)'=2x-2=2(x-1)\] Now set y'=0

myininaya (myininaya):

and solve for x

myininaya (myininaya):

Like this: 2(x-1)=0

myininaya (myininaya):

But guess what happens?

myininaya (myininaya):

x=1 is not btw -2 and 0 so we do not care about x=1 We only have to look the endpoints

OpenStudy (anonymous):

ok

myininaya (myininaya):

So evaluate y for x=0 Then evaluate y for x=-2

myininaya (myininaya):

Which is higher output and which is the lower output?

OpenStudy (anonymous):

how do i evaluate?

myininaya (myininaya):

\[y|_{x=0}=0^2-2(0)-7=-7\] \[y|_{x=-2}=(-2)^2-2(-2)-7=4+4-7=8-7=1\]

myininaya (myininaya):

See I evaluated y for x=0 and x=-2

myininaya (myininaya):

Now 1 is higher than -7 So 1 is the max value and -7 is the min value on the interval [-2,0]

OpenStudy (anonymous):

if so why do i need to derive the calculus? if i can just substitute into the original function to get the max and the mini?

myininaya (myininaya):

Because we needed to figure out if we had any critical numbers btw the endpoints

myininaya (myininaya):

For example if the interval was [0,2] we would have gotten something differently ....

myininaya (myininaya):

We had x=1 is the critical number 1 is btw 0 and 2 so we have to check both endpoints and the critical number for this one

OpenStudy (anonymous):

wat of in cases like this Find the maximum and minimum values of the function on the given interval. y = (3 - x)/(x^2 + 9 x), text( ) text( ) [3, 10]

myininaya (myininaya):

Find y' Set y'=0 Also find where y' DNE These will give you your critical numbers Make sure you only look at the critical numbers btw 3 and 10

OpenStudy (anonymous):

u can just go on with ur example,if i learn i can just solve this one i posted and the rest. as u were saying if 1 happens to be with the given intervals,how will i get the max and the mini from there?

myininaya (myininaya):

\[y'=\frac{-1(x^2+9x)-(3-x)(2x+9)}{(x^2+9x)^2}=\frac{-x^2-9x-(6x+27-2x^2-9x)}{(x^2+9x)^2}\] \[=\frac{-x^2-9x-(-2x^2-3x+27)}{(x^2+9x)^2}\] =\[\frac{-x^2-9x+2x^2+3x-27}{(x^2+9x)^2}=\frac{x^2-6x-27}{(x^2+9x)^2}\] So where is y'=0 and where does y' DNE ?

myininaya (myininaya):

Oh you want to do the one I was doing lol sorry

OpenStudy (anonymous):

y does not exist when x=0

OpenStudy (anonymous):

dont worry just go on with this

OpenStudy (anonymous):

dis will give me a better understanding

myininaya (myininaya):

Ok so we had \[y=x^2-2x-7, [0,2]\] So we find y' just like we did before y'=2x-2=2(x-1) Set y'=0 and we get x=1 So 1 is btw the endpoints so we have to check all three of these x-values

myininaya (myininaya):

\[y|_{x=0}=0^2-2(0)-7=-7\] \[y|_{x=1}=1^2-2(1)-7=1-2-7=1-9=-8\] \[y|_{x=2}=2^2-2(2)-7=-7\] -8 is smallest so -8 is the min value -7 is highest so -7 is the max value on [0,2]

myininaya (myininaya):

Ok so if we were looking at this new problem you put up

OpenStudy (anonymous):

yeah

myininaya (myininaya):

\[y'=\frac{x^2-6x-27}{(x^2+9x)^2}\]

myininaya (myininaya):

To find where y'=0 Set top=0 and solve for x \[x^2-6x-27=0 => (x-9)(x+3)=0 \] x=? or x=? ----------------------------- To find where y' DNE Set bottom=0 and solve for x \[(x^2+9x)^2=0 => x^2+9x=0 =>x(x+9)=0\] x=? or x=? -------------------------------

myininaya (myininaya):

What are the critical numbers?

OpenStudy (anonymous):

-9 and 0?

myininaya (myininaya):

and ...

myininaya (myininaya):

9 and -3 right?

myininaya (myininaya):

now we only want to look at the critical numbers on the interval given

myininaya (myininaya):

So which are those?

OpenStudy (anonymous):

why -3?

myininaya (myininaya):

-3+3=0

myininaya (myininaya):

You did solve both equations right?

OpenStudy (anonymous):

i solved for x(x+9)=0

myininaya (myininaya):

What about the other equation I had there to solve?

myininaya (myininaya):

Did you read everything I typed?

OpenStudy (anonymous):

oh! now i see the equation,am sorry :)

OpenStudy (anonymous):

u can go on

myininaya (myininaya):

ok so the interval was [3,10] right?

OpenStudy (anonymous):

yep

myininaya (myininaya):

Now which of the critical numbers we found are btw 3 and 10?

OpenStudy (anonymous):

9

myininaya (myininaya):

ok so we need to evaluate y for x=3 evaluate y for x=9 evaluate y for x=10

myininaya (myininaya):

\[y=\frac{3-x}{x^2+9x}\] \[y|_{x=3}=\frac{3-3}{3^2+9(3)}=0\] \[y|_{x=9}=\frac{3-9}{9^2+9(9)}=\frac{-6}{2(81)}=\frac{-3}{81}\] \[y|_{x=10}=\frac{3-10}{10^2+9(10)}=\frac{-7}{190}\]

myininaya (myininaya):

Which is the lowest output?

OpenStudy (anonymous):

-7/190

OpenStudy (anonymous):

so that's my mini right?

myininaya (myininaya):

-3/81 is smallest so -3/81 is the min value 0 is highest so 0 is the max value

OpenStudy (anonymous):

ok,so i can use this procedure to solve the rest?

myininaya (myininaya):

what do you mean?

OpenStudy (anonymous):

like a question like this y = (sqrt(5 + x^2)) - 4 x, [5, 6] i can use ur procedure to determine the max and the mini?

myininaya (myininaya):

of course y'=0 and y' DNE Use all critical numbers btw the endpoints given and the endpoints and plug them into y see what the lowest output (min value) is and then see what the highest output (max value) is

OpenStudy (anonymous):

ok thanks alot for your help :)

myininaya (myininaya):

By the way when finding values for y' DNE make sure that y exists for those x values

OpenStudy (anonymous):

oh ok sure :)

OpenStudy (anonymous):

@myininaya i have a problem with the remaining two quesrtions

myininaya (myininaya):

two?

OpenStudy (anonymous):

yeah in this situation,i set my function=0 and got x-2((4+x^2)^1/2)

OpenStudy (anonymous):

is my cn just 0 in this case?

myininaya (myininaya):

I don't understand...Are these the same type of questions as before?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

So what is y?

OpenStudy (anonymous):

y = (sqrt(4 + x^2)) - 2 x, [4, 5]

myininaya (myininaya):

\[y'=\frac{x}{\sqrt{4+x^2}}-2\]Is that what you got for y'?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the i set it to 0

myininaya (myininaya):

Yes

myininaya (myininaya):

\[\frac{x-2\sqrt{4+x^2}}{\sqrt{4+x^2}}=0\]

OpenStudy (anonymous):

got x-2(4+x^2)^1/2

OpenStudy (anonymous):

yea

myininaya (myininaya):

\[x-2 \sqrt{4+x^2}=0 \] \[x=2 \sqrt{4+x^2} => x^2=4(4+x^2) \text{ squaring both sides}\]

OpenStudy (anonymous):

ok,so my cn will be?

myininaya (myininaya):

\[x^2=16+4x^2 =. x^2-4x^2=16 => -3x^2=16 => x^2=\frac{16}{-3}\] But this will give us imaginary x So we don't have any x-values that satisfy y'=0 So now look for when y' DNE

myininaya (myininaya):

But again no such values exist

myininaya (myininaya):

So last thing to do is just plug in endpoints

OpenStudy (anonymous):

which i s [4 and 5] ?

myininaya (myininaya):

\[y|_{x=4} =\sqrt{4+4^2}-2(4) \] \[y|_{x=5}=\sqrt{4+5^2}-2(5)\]

OpenStudy (anonymous):

so that will be my max and mini?

myininaya (myininaya):

the lowest ouput=min value the highest output=max value

myininaya (myininaya):

\[y=5e^x-e^{3x}\] ?

OpenStudy (anonymous):

yeah exactly {-1,1}